OpenStudy (anonymous):

How would you write 0.3y + 1.2 = –0.4x in standard form? How would you write 0.3y + 1.2 = –0.4x in standard form? @Mathematics

OpenStudy (anonymous):

Standard form is y=mx+b So all we have to do is put the y on one side of the equal sign, alone. 0.3y+1.2=-0.4x *subtract 1.2* .3y=-0.4x-1.2 *divide .3y by .3, to get the y alone* y= -(.4x/.3) -(1.2/.3) *simplify* y= -.4x/.3 - 4 ((because -1.2/.3 = -4)

OpenStudy (anonymous):

Or you could write the final answer as y= -.4/.3 x - 4 if you want it to look even more like standard form. Although it's really the same thing, and you should get credit for either. It's only a very subtle insignificant difference.